# https://www.mitre.org/

### v1

@@ Abstract
We explore the scaling behaviors of artificial intelligence to establish practical techniques for training foundation models on high-resolution electro-optical (EO) datasets that exceed the current state-of-the-art scale by orders of magnitude. Modern multimodal machine learning (ML) applications, such as generative artificial intelligence (GenAI) systems for image captioning, search, and reasoning, depend on robust, domain-specialized encoders for non-text modalities. In natural-image domains where internet-scale data is plentiful, well-established scaling laws help optimize the joint scaling of model capacity, training compute, and dataset size. Unfortunately these relationships are much less well-understood in high-value domains like remote sensing (RS). Using over a quadrillion pixels of commercial satellite EO data and MITRE’s Federal AI Sandbox, we train progressively larger vision transformer (ViT) backbones, report successes and failure modes observed at petascale, and analyze implications for bridging domain gaps across additional RS modalities. We observe that even at this scale, performance is consistent with a data-limited regime rather than a model parameter-limited one. These practical insights are intended to inform data-collection strategies, compute budgets, and optimization schedules that advance the future development of frontier-scale RS foundation models.
@@ 1 Introduction
Recent advances in machine learning (ML) have been propelled by the development of large-scale models and foundation architectures, particularly in the context of generative artificial intelligence (GenAI) and multi-modal applications. Central to the success of these systems are performant data encoders which transform raw data into representations useful in downstream tasks. While scaling laws and architectural best practices for vision transformers (ViTs) have been extensively studied in domains such as natural imagery [scalingvits] [nlpscaling], their applicability to specialized domains like remote sensing (RS) remains underexplored.
Remote sensing data, especially from electro-optical (EO) commercial satellites, presents unique challenges and opportunities for ML. Unlike natural images, RS data is characterized by diverse spectral, spatial, and temporal properties, as well as limited availability of labeled, in-domain training samples. This scarcity has hindered the development and scaling of foundation models (FMs) tailored to RS, leaving several fundamental questions unanswered. Notably, the relationship between training dataset size and model performance at scale for in-domain RS FMs is poorly understood.
Addressing these gaps is critical for advancing the state of the art in RS ML applications, which underpin a wide range of societal and scientific needs, from environmental monitoring to urban planning. To this end, we leverage an unprecedented scale of EO satellite imagery totaling over a peta-pixel and employ multi-task pretraining (MTP) [mtp] and Scale-Aware Masked Autoencoder (ScaleMAE) [scalemae] foundation model architectures. By integrating supervised optimization with open-source geographic information system (GIS)-derived labels, we are able to construct large, high-quality training datasets. Furthermore, by utilizing large-scale graphics processing units (GPUs) resources, specifically an NVIDIA DGX H100 SuperPOD in MITRE’s Federal AI Sandbox, we systematically scale ViT training regimes to assess whether modern frontier RS foundation models remain constrained by data availability.
In this work, we empirically characterize the scaling behaviors of ViT-based RS FMs, evaluating their performance and analyzing their capacity to address domain gaps across additional RS modalities. Our findings provide new insights into the design and training of RS FMs, with implications for both the ML and RS communities.
@@ 1.1 Guiding Questions
This study is organized around four guiding questions central to scaling RS encoders: (a) under what conditions increasing batch size improves training time or convergence; (b) which learning-rate schedules remain stable for continual pre-training of RS models at scale; (c) how performance scales with in-domain dataset size and the rate at which gains taper; and (d) how predictably performance improves when scaling ViT model parameters. These questions structure our experimental design and analyses to develop practical strategies for future RS FMs training.
@@ 1.3 Contributions
  • Empirical model-scaling laws: across 86M-1.8B parameters we see effectively flat returns indicating that model scaling alone provides negligible benefit under typical data size regimes.
@@ 2.1 Foundation Models and Scaling Laws in Vision
The adoption of transformer units [attention] [vits] revolutionized representation learning by introducing self-attention mechanisms that scale efficiently with data and compute. This paradigm shift enabled the rise of foundation models, large, pre-trained systems capable of generalizing across diverse downstream tasks with minimal adaptation. Studies on neural network scaling later showed that model performance frequently follows power-law relationships with respect to model size, dataset size, and compute. Previous research [nlpscaling] demonstrated these trends for transformer-based language models showing that loss scales predictably and that compute can be optimally distributed across parameters, data, and training steps. In vision, research [scalingvits] extended scaling analysis to vision transformers (ViTs), finding a compute performance pareto frontier and double-saturating power law indicating that data remains the limiting factor even for billion-scale image datasets. Larger models continue to benefit from more data, while smaller ones quickly saturate. Scaling-law methods were quickly refined through improved curve fitting [revistneuralscaling], enhancing extrapolation across new scales and modalities and enabling more accurate prediction of compute efficiency. Together these works established scaling laws as a predictive framework for large-scale model design. Our study extends this foundation to geospatial foundation models, where domain heterogeneity, sensor diversity, and spatial temporal structure may yield distinct scaling dynamics.
@@ 2.2 Vision for Remote Sensing
To address the diversity and scale of earth observation data, researchers have begun to assemble increasingly larger, more heterogenous datasets and corresponding pretraining strategies. SatlasPretrain [satlaspretrain] created one of the largest labeled RS pretraining dataset to date with multiple sensors and tasks: $1\times 10^{13}$ pixels over 21 million $\mathrm{km}^{2}$ with 137 distinct categories. The model demonstrated that large-scale joint pretraining improves downstream performance on a variety of segmentation, object detection, and change detection tasks marking an early step toward foundation-scale generalization in RS. Similarly, Prithvi [prithvi] explored self-supervised pretraining for multispectral imagery at continental scale (4.2M globally distributed samples), revealing that transformer models can learn general geospatial representations transferable across multiple datasets.
Architecturally, recent work has been focused on incorporating scale awareness and geospatial priors into vision transformers. ScaleMAE [scalemae] introduced a scale-aware masked autoencoder that anchors positional encoding to real-world ground sample distance (GSD), enabling consistent multiresolution learning across diverse imagery sources. Other work proposed Multi-Task Pretraining (MTP) [mtp] frameworks that unify segmentation, detection, and classification objectives, while SkySense [skysense] extended foundational pretraining to billions of parameters and multimodal inputs, both temporal and contextual, reflecting the domains’ upward progress. Efforts that have explicitly expanded parameter scaling, training transformer backbones ranging from 86M to 2B parameters [billionscale] observed consistent performance in object detection and segmentation regardless of backbone size, likely due to a lack of commensurate data scaling with parameter scaling, as they leveraged the same 1M examples as the original MTP effort [mtp]. Meanwhile, DINOv3 [dinov3], provides evidence that scaling model and data size in self-supervision can produce dense transferable representations that generalize to RS domains, starting both from natural and RS images, with corpuses of 1.7B images and 493M images, respectively. The increase in RS dataset sizes across the field is summarized in table 1.
@@ 2.3 Summary and Gaps
Overall, these developments highlight a clear trend: remote sensing has entered a new scale of vision FM development. With increasing dataset size, parameter counts, and compute budgets, the field’s next step is to quantitatively characterize scaling laws to determine how performance scales within the unique characteristics of geospatial imagery. Following the success of DINOv3 [dinov3] in creating relevant RS representations, the field would benefit from precise quantification of training tradeoffs to allow for right-sizing of datasets, model architectures, and compute budgets for specific domain needs. Most prior work emphasizes best-case performance and model design, while comparatively little documents the operational regimes where scaling stalls, destabilizes, or provides marginal returns in RS. Our study complements these advances by quantifying when and how scaling pays off in RS, and by reporting the failure modes and lessons learned that turn scaling laws into practical training practice.
@@ 3 Dataset
We constructed Akupara, a global remote sensing dataset comprising over a peta-pixel ($1\times 10^{15}$) of high-resolution Maxar satellite imagery, the largest remote sensing dataset to date for FM training. The “Akupara-1M” collection contains over 1 million full satellite images (equivalent to $\sim 4$ billion $512\times 512$ chips) with 415 million $\mathrm{km}^{2}$ of gross coverage (excluding USA due to collection constraints), with both single-channel panchromatic (PAN) imagery and eight-band multi-spectral imagery (MSI), and a ground sample distance (GSD) range of 0.3 meters to 1.2 meters. Images were collected primarily from from 2020-2024 with <2% of imagery from 2007-2020. Table 1 shows a comparison of Akupara to other large remote sensing datasets.
@@ 3.1 Label Generation
We developed the automated Geospatial Data Augmentation (G-DAUG) pipeline, a scalable geospatial labeling framework based on the SatlasPretrain [satlaspretrain] labeling method for large-scale weak supervision and automated annotation. Each $1024\times 1024$ image chip is paired with OpenStreetMap (OSM) vector data retrieved via PostGIS queries within the chip’s geographic footprint. The pipeline converts chip pixel coordinates to geographic bounds using metadata from the source orthorectified imagery, then filters OSM geometries by a curated ontology of 37 object classes (based on empirical findings in [satlaspretrain] that reveal categories with higher fidelity labels). Geometries are reprojected to the image coordinate reference system, clipped to the chip extent, and then rasterized into high-resolution binary segmentation masks. We produce instance segmentation masks, semantic segmentation masks, and rotated bounding boxes labels (as shown in figure 2) required for supervised training, such as with the Multi-Task Pretraining paradigm. This pipeline provides a reproducible, automated system for generating consistent, chip-aligned supervision across global-scale imagery datasets, bridging open vector sources and high-resolution raster training data.
Figure 2: The Geospatial Data Augmentation (G-DAUG) pipeline leverages OpenStreetMap data and segmentation algorithms to automatically create geospatial data labels. We create pixelwise segmentation masks, as shown, which are also converted bounding boxes and instance segmentation masks for additional supervisory tasks.
@@ 4.1 Model Architectures
We leverage two different model configurations for our experiments: Multi-Task Pretraining (MTP), which is fully supervised and uses multi-task optimization across multiple label types; and Scale-Aware Masked Autoencoder (ScaleMAE), which is based on self-supervision and scale invariance.
Multi-Task Pretraining (MTP) is a training paradigm that consists of a single backbone (or shared encoder) and a set of task-specific decoders [mtp]. During pretraining, the model is optimized jointly on three remote sensing tasks: instance segmentation, semantic segmentation, and rotated object detection, all of which are supervised with mask and bounding box labels, as appropriate for each task. The shared encoder produces a multi-scale feature pyramid that serves as input to each of the task-specific decoders. This framework is targeted towards large foundation models, supporting both CNN and ViT architectures. Additionally, when a ViT is used, the traditional multi-head full attention is replaced by Rotated Varied-Size Window Attention (RVSA) which helps to account for the various orientations of RS objects by rotating and rescaling local windows, allowing the transformer to attend to objects at arbitrary orientations and scales. After pretraining is completed, the model can be finetuned on a downstream RS task.
Scale-Aware Masked Autoencoder (ScaleMAE) is a self-supervised pretraining framework designed for multiscale geospatial imagery [scalemae]. It extends the standard Masked Autoencoder (MAE) by encoding real-world scale information, such as ground sample distance (GSD), directly into the positional embeddings to enable consistent representations across sensors and resolutions. The architecture uses a ViT encoder and lightweight decoder trained to reconstruct masked image patches, enforcing scale-invariant feature learning through reconstruction losses measured relative to physical, rather than pixel, scale. This approach allows the model to learn semantically aligned features across heterogenous imagery without requiring resampling or manual normalization. Once pretrained, the encoder can be finetuned for downstream RS tasks.
@@ 4.2 Balancing Training Time and Compute Efficiency
To evaluate the relationship between batch size, training time, and compute efficiency, we conducted a series of critical batch size experiments using ScaleMAE ViT-Base architecture [scalemae]. This smaller model configuration was chosen to enable multiple experimental repetitions across varying batch sizes and random seeds while maintaining computational feasibility. Each run used the same dataset, hyperparameters, and optimizer (AdamW), ensuring consistent optimization setting across runs. During these runs training was performed on a single NVIDIA DGX node containing eight H100 GPUs connected by NVLink, progressing through early to mid-stage training before hitting a target validation loss of 0.07 mean squared error (MSE) to capture the regime where batch scaling effects are most prominent. Batch sizes selected were 128, 256, 384, 512 and results were averaged over multiple seeds to mitigate variance (see figure 5). These settings isolate our ability to manipulate training throughput without confounding optimizer dynamics, enabling a direct read on whether batch scaling accelerates convergence. We report the outcomes and implications for compute budgeting in Section 5.1.
4.3 Optimal Learning Rate (LR) Scheduling
To determine ranges of functional learning rates for large-scale geospatial pretraining, we conducted a controlled sweep across multiple rates for MTP-trained models. Models were trained with the same dataset, batch size and Warmup, Stable, Decay (WSD) scheduling strategy with stabilized checkpointing to isolate the effect of the base learning rate [wsd]. Following best practices for responsible compute allocation, we adopt a fail fast protocol to triage candidate base learning rates before committing full WSD runs. Concretely, we evaluate a broader set of LR candidates through 1.6k warmup iterations with Akupara-5k, under identical data, batch, and optimizer settings on four DGX nodes (32 H100s). The rates used during this preliminary test are shown in figure 3.
@@ 4.3 Optimal Learning Rate (LR) Scheduling
Figure 4: We train three selected stable learning rates with the Warmup, Stable, Decay (WSD) schedule on Akupara-10k, comprising a 10% warmup, an 80% stable plateau at the target rate, and an exponential decay initiated when approximately 10% of the data remains. The curves trace the full WSD cycle for each rate, illustrating the transition from warmup to plateau (listed stable LR) and subsequent decay.

### v2

@@ Abstract
We explore the scaling behaviors of artificial intelligence to establish practical techniques for training foundation models on high-resolution electro-optical (EO) datasets that exceed the current state-of-the-art scale by orders of magnitude. Modern multimodal machine learning (ML) applications, such as generative artificial intelligence (GenAI) systems for image captioning, search, and reasoning, depend on robust, domain-specialized encoders for non-text modalities. In natural image domains where internet-scale data is plentiful, well established scaling laws help optimize the joint scaling of model capacity, training compute, and dataset size. Unfortunately these relationships are much less well understood in high-value domains like remote sensing (RS). Using over a quadrillion pixels of commercial satellite EO data and MITRE’s Federal AI Sandbox, we train progressively larger vision transformer (ViT) backbones, report successes and failure modes observed at petascale, and analyze implications for bridging domain gaps across additional RS modalities. We observe that even at this scale, performance is consistent with a data-limited regime rather than a model parameter-limited one. These practical insights are intended to inform data collection strategies, compute budgets, and optimization schedules that advance the future development of frontier scale RS foundation models.
@@ 1 Introduction
Recent advances in ML have been propelled by the development of large-scale models and foundation architectures, particularly in the context of GenAI and multimodal applications. Central to the success of these systems are performant data encoders which transform raw data into representations useful in downstream tasks. While scaling laws and architectural best practices for ViTs have been extensively studied in domains such as natural imagery [14] [6], their applicability to specialized domains like RS remains underexplored.
Remote sensing data, especially from EO commercial satellites, presents unique challenges and opportunities for ML. Unlike natural images, RS data is characterized by diverse spectral, spatial, and temporal properties, as well as limited availability of labeled, in-domain training samples. This scarcity has hindered the development and scaling of foundation models (FMs) tailored to RS, leaving several fundamental questions unanswered. Notably, the relationship between training dataset size and model performance at scale for in-domain RS FMs is poorly understood.
Addressing these gaps is critical for advancing the state of the art in RS ML applications, which underpin a wide range of societal and scientific needs, from environmental monitoring to urban planning. To this end, we leverage an unprecedented scale of EO satellite imagery totaling over a peta-pixel and employ multi-task pretraining (MTP) [12] and scale-aware masked autoencoder (ScaleMAE) [7] foundation model architectures. By integrating supervised optimization with open-source geographic information system (GIS) derived labels, we are able to construct large, high-quality training datasets. Furthermore, by utilizing large-scale graphics processing units (GPUs) resources, specifically an NVIDIA DGX H100 SuperPOD in MITRE’s Federal AI Sandbox, we systematically scale ViT training regimes to assess whether modern frontier RS foundation models remain constrained by data availability.
In this work, we empirically characterize the scaling behaviors of ViT-based RS FMs, evaluating their performance and analyzing their capacity to address domain gaps across additional geospatial modalities. Our findings provide new insights into the design and training of remote sensing FMs, with implications for both the ML and RS communities.
@@ 1.1 Guiding Questions
This study is organized around four guiding questions central to scaling RS encoders: (a) under what conditions increasing batch size improves training time or convergence; (b) which learning rate (LR) schedules remain stable for continual pretraining of RS models at scale; (c) how performance scales with in-domain dataset size and the rate at which gains taper; and (d) how predictably performance improves when scaling ViT model parameters. These questions structure our experimental design and analyses to develop practical strategies for future RS FMs training.
@@ 1.3 Contributions
  • Empirical model scaling laws: across 86M-1.8B parameters we see effectively flat returns indicating that model scaling alone provides negligible benefit under typical data size regimes.
@@ 2.1 Foundation Models and Scaling Laws in Vision
The adoption of transformer units [11] [4] revolutionized representation learning by introducing self-attention mechanisms that scale efficiently with data and compute. This paradigm shift enabled the rise of foundation models, large, pretrained systems capable of generalizing across diverse downstream tasks with minimal adaptation. Studies on neural network scaling later showed that model performance frequently follows power-law relationships with respect to model size, dataset size, and compute. Previous research [6] demonstrated these trends for transformer-based language models showing that loss scales predictably and that compute can be optimally distributed across parameters, data, and training steps. In vision, research [14] extended scaling analysis to ViTs, finding a compute performance pareto frontier and double-saturating power law indicating that data remains the limiting factor even for billion-scale image datasets. Larger models continue to benefit from more data, while smaller ones quickly saturate. Scaling-law methods were quickly refined through improved curve fitting [1], enhancing extrapolation across new scales and modalities and enabling more accurate prediction of compute efficiency. Together these works established scaling laws as a predictive framework for large-scale model design. Our study extends this foundation to geospatial foundation models, where domain heterogeneity, sensor diversity, and spatial temporal structure may yield distinct scaling dynamics.
@@ 2.2 Vision for Remote Sensing
To address the diversity and scale of earth observation data, researchers have begun to assemble increasingly larger, more heterogenous datasets and corresponding pretraining strategies. SatlasPretrain [2] created one of the largest labeled RS pretraining datasets to date with multiple sensors and tasks: $1\times 10^{13}$ pixels over 21 million $\mathrm{km}^{2}$ with 137 distinct categories. The model demonstrated that large-scale joint pretraining improves downstream performance on a variety of segmentation, object detection, and change detection tasks marking an early step toward foundation-scale generalization in RS. Similarly, Prithvi [9] explored self-supervised pretraining for multispectral imagery at continental scale (4.2M globally distributed samples), revealing that transformer models can learn general geospatial representations transferable across multiple datasets.
Architecturally, recent work has been focused on incorporating scale awareness and geospatial priors into vision transformers. ScaleMAE [7] introduced a scale-aware masked autoencoder that anchors positional encoding to real-world ground sample distance (GSD), enabling consistent multiresolution learning across diverse imagery sources. Other work proposed Multi-Task Pretraining (MTP) [12] frameworks that unify segmentation, detection, and classification objectives, while SkySense [5] extended foundational pretraining to billions of parameters and multimodal inputs, both temporal and contextual, reflecting the domains’ upward progress. Efforts that have explicitly expanded parameter scaling, training transformer backbones ranging from 86M to 2B parameters [3], observed consistent performance in object detection and segmentation regardless of backbone size, likely due to a lack of commensurate data scaling with parameter scaling, as they leveraged the same 1M examples as the original MTP effort [12]. Meanwhile, DINOv3 [8], provides evidence that scaling model and data size in self-supervision can produce dense transferable representations that generalize to RS domains, starting both from natural and RS images, with corpuses of 1.7B image chips and 493M image chips, respectively. The increase in RS dataset sizes across the field is summarized in table 1.
@@ 2.3 Summary and Gaps
Overall, these developments highlight a clear trend: remote sensing has entered a new scale of vision FM development. With increasing dataset size, parameter counts, and compute budgets, our next step is to quantitatively characterize scaling laws to determine how performance scales within the unique characteristics of geospatial imagery. Following the success of DINOv3 [8] in creating relevant RS representations, the field would benefit from precise quantification of training tradeoffs to allow for right-sizing of datasets, model architectures, and compute budgets for specific domain needs. Most prior work emphasizes best-case performance and model design, while comparatively little documents the operational regimes where scaling stalls, destabilizes, or provides marginal returns in RS. Our study complements these advances by quantifying when and how scaling pays off in RS, and by reporting the failure modes and lessons learned that turn scaling laws into practical training practice.
@@ 3 Dataset
We constructed Akupara, a global remote sensing dataset comprising over a peta-pixel ($1\times 10^{15}$) of high-resolution Maxar satellite imagery, the largest remote sensing dataset to date for FM training. The “Akupara-1M” collection contains over 1 million full satellite images (equivalent to $\sim 4$ billion $512\times 512$ chips) with 415 million $\mathrm{km}^{2}$ of gross coverage (excluding USA due to collection constraints), with both single-channel panchromatic (PAN) imagery and eight-band multi-spectral imagery (MSI), and a ground sample distance (ground sampling distance (GSD)) range of 0.3 meters to 1.2 meters. Images were collected primarily from from 2020-2024 with <2% of imagery from 2007-2020. Table 1 shows a comparison of Akupara to other large remote sensing datasets.
@@ 3.1 Label Generation
We developed the automated G-DAUG pipeline, a scalable geospatial labeling framework based on the SatlasPretrain [2] labeling method for large-scale weak supervision and automated annotation. Each $1024\times 1024$ image chip is paired with OpenStreetMap (OSM) vector data retrieved via PostGIS queries within the chip’s geographic footprint. The pipeline converts chip pixel coordinates to geographic bounds using metadata from the source orthorectified imagery, then filters OSM geometries by a curated ontology of 37 object classes (based on empirical findings in [2] that reveal categories with higher fidelity labels). Geometries are reprojected to the image coordinate reference system, clipped to the chip extent, and then rasterized into high-resolution binary segmentation masks. We produce instance segmentation masks, semantic segmentation masks, and rotated bounding boxes labels (as shown in figure 2) required for supervised training, such as with the Multi-Task Pretraining paradigm. This pipeline provides a reproducible, automated system for generating consistent, chip-aligned supervision across global-scale imagery datasets, bridging open vector sources and high-resolution raster training data.
Figure 2: The G-DAUG pipeline leverages OSM data and segmentation algorithms to automatically create geospatial data labels. We create pixelwise segmentation masks, as shown, which are also converted bounding boxes and instance segmentation masks for additional supervisory tasks.
@@ 4.1 Model Architectures
We leverage two different model configurations for our experiments: MTP, which is fully supervised and uses multi-task optimization across multiple label types; and ScaleMAE, which is based on self-supervision and scale invariance.
MTP is a training paradigm that consists of a single backbone (or shared encoder) and a set of task-specific decoders [12]. During pretraining, the model is optimized jointly on three remote sensing tasks: instance segmentation, semantic segmentation, and rotated object detection, all of which are supervised with mask and bounding box labels, as appropriate for each task. The shared encoder produces a multi-scale feature pyramid that serves as input to each of the task-specific decoders. This framework is targeted towards large foundation models, supporting both Convolutional Neural Network (CNN) and ViT architectures. Additionally, when a ViT is used, the traditional multi-head full attention is replaced by rotated varied-size attention (RVSA) which helps to account for the various orientations of RS objects by rotating and rescaling local windows, allowing the transformer to attend to objects at arbitrary orientations and scales. After pretraining is completed, the model can be finetuned on a downstream RS task.
ScaleMAE is a self-supervised pretraining framework designed for multiscale geospatial imagery [7]. It extends the standard masked autoencoder (MAE) by encoding real-world scale information, such as GSD, directly into the positional embeddings to enable consistent representations across sensors and resolutions. The architecture uses a ViT encoder and lightweight decoder trained to reconstruct masked image patches, enforcing scale-invariant feature learning through reconstruction losses measured relative to physical, rather than pixel, scale. This approach allows the model to learn semantically aligned features across heterogenous imagery without requiring resampling or manual normalization. Once pretrained, the encoder can be finetuned for downstream RS tasks.
@@ 4.2 Balancing Training Time and Compute Efficiency
To evaluate the relationship between batch size, training time, and compute efficiency, we conducted a series of critical batch size experiments using ScaleMAE ViT-Base architecture [7]. This smaller model configuration was chosen to enable multiple experimental repetitions across varying batch sizes and random seeds while maintaining computational feasibility. Each run used the same dataset, hyperparameters, and optimizer (AdamW), ensuring consistent optimization setting across runs. During these runs training was performed on a single NVIDIA DGX node containing eight H100 GPUs connected by NVLink, progressing through early to mid-stage training before hitting a target validation loss of 0.07 mean squared error (MSE) to capture the regime where batch scaling effects are most prominent. Batch sizes selected were 128, 256, 384, 512 and results were averaged over multiple seeds to mitigate variance (see figure 5). These settings isolate our ability to manipulate training throughput without confounding optimizer dynamics, enabling a direct read on whether batch scaling accelerates convergence. We report the outcomes and implications for compute budgeting in Section 5.1.
4.3 Optimal LR Scheduling
To determine ranges of functional learning rates for large-scale geospatial pretraining, we conducted a controlled sweep across multiple rates for MTP-trained models. Models were trained with the same dataset, batch size and WSD scheduling strategy with stabilized checkpointing to isolate the effect of the base learning rate [13]. Following best practices for responsible compute allocation, we adopt a fail fast protocol to triage candidate base learning rates before committing full WSD runs. Concretely, we evaluate a broader set of LR candidates through 1.6k warmup iterations with Akupara-5k, under identical data, batch, and optimizer settings on four DGX nodes (32 H100s). The rates used during this preliminary test are shown in figure 3.
@@ 4.3 Optimal LR Scheduling
Figure 4: We train three selected stable learning rates with the WSD schedule on Akupara-10k, comprising a 10% warmup, an 80% stable plateau at the target rate, and an exponential decay initiated when approximately 10% of the data remains. The curves trace the full WSD cycle for each rate, illustrating the transition from warmup to plateau (listed stable LR) and subsequent decay.